OpenStudy (anonymous):

A motorcycle is depreciating at 12% per year, every year. A student's $11,250 motorcycle depreciating at this rate can be modeled by the equation V(t) = 11,250(0.88)t. What is an equivalent equation for this vehicle as a monthly depreciation and, using this equation, what is the motorcycle worth (rounded to the nearest hundred dollar) 8 years after purchase?

OpenStudy (anonymous):

V(t) = 11,250(0.9894)12t, $4,000 V(t) = 11,250(0.8800)12t, $10,300 V(t) = 11,250(1.12)−t, $4,500 V(t) = 11,250(0.9894)t, $4,000

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

the 12t is correct because there are 12 months in 1 year. t will now be the number of months. But I cant see how they get 0.9894.

OpenStudy (anonymous):

thats the same thing i said but ill go with the first one , thanks